OpenStudy (rjtrain1):

Reduce 36/45 to lowest terms and write the denominator in the blank will medal and fan

OpenStudy (rjtrain1):

no specific answers

satellite73 (satellite73):

what does "no specific answers" mean?

OpenStudy (rjtrain1):

like any answer is ok not multiple choice

satellite73 (satellite73):

if your multiplication memory is good, you will recognize that both 36 and 45 are multiples of some number

OpenStudy (rjtrain1):

what

satellite73 (satellite73):

can you think of a number that goes in to 36 and also goes in to 45?

OpenStudy (tkhunny):

Hint: "No Specific Answer" means we are not given the final answer and we must figure it out for ourselves. This is good. Never look at that, even if it is given.

OpenStudy (rjtrain1):

6*6=36 and8*5=45

satellite73 (satellite73):

actually \(8\times 5=40\)

satellite73 (satellite73):

what i was asking was, can you think of ONE number that goes in to both 36 and 45

OpenStudy (mantar0078):

Once you find the number that goes into both 36 and 45 you divide that number into 36 and 45 to get your new fraction. If the fraction is something like 5/10 then you would simplify that again to get 1/2, which can't be simplified.

OpenStudy (mantar0078):

For instance 5 is a multiple of 15 and 20 so they would be simplified to 3/4.

OpenStudy (amazingme_16):

Well, since both of these numbers have 9 as the highest common denominator you could divide 36 by 9 and get 4, and 45 divided by 9 and get 5. Now that you have divided you would match the the numbers with what you divided it by. For instance, you would get 4/5.

OpenStudy (amazingme_16):

Sorry meant highest common factor

OpenStudy (rjtrain1):

wait 9 does

OpenStudy (osprey):

36/45 to lowest terms 12/15 4/5 keep dividing top and bottom by 3

OpenStudy (osprey):

http://perendis.webs.com

OpenStudy (rjtrain1):

and I like the exzact answer please

OpenStudy (rjtrain1):

no advertising sites

OpenStudy (mantar0078):

We can't just give you the answer. OpenStudy will suspend our account if we do. You have to do at least something to figure out the answer yourself. Figuring out which number goes into 36 and 45 really isn't that hard.